OpenStudy (jennychan12):

Find the area bounded by the curves y = x/2, y = sqrt{x} , and x = 9.

OpenStudy (jennychan12):

\[y = \frac{ 1 }{ 2 }x \] \[y = \sqrt{x}\] \[x = 9\]  i assume it's the shaded region. so i know it's \[\int\limits_{0}^{9}\] but i don't know if the inside is sqrt(x) - x/2 or if its x/2 - sqrt(x) plus the dx of course.